Suppose product price is fixed at $24; MR = MC at Q = 200; AFC = $6; AVC = $16. What do you advise this firm to do?

a.

Decrease output.

b.

Stay at the current output; the firm is losing $200.

c.

Increase output.

d.

Stay at the current output; the firm is earning a profit of $400.
Answer #1

Answer

Option d

Stay at the current output; the firm is earning a profit of $400

ATC=AFC+AVC=6+16=22

P=24

Q=200

Profit=(P-ATC)*Q=(24-22)*200=$400

the firm makes a profit of $400